A hydraulic press consists of two cylinders of cross-sectional area 0.2m and 5.0m. The piston in there smaller cylinder is pushed down with a force of 100N through a distance of 0.2m. Calculate the force exerted by the piston in the larger cylinder.
1
Expert's answer
2018-09-11T12:49:08-0400
We can find the force exerted by the piston in the larger cylinder from the hydraulic press formula:
F_1/A_1 =F_2/A_2 ,
here, F_1=100 N is the force exerted by the piston in the smaller cylinder, F_2 is the force exerted by the piston in the larger cylinder, A_1=0.2 m^2 is the cross-sectional area of the smaller cylinder, A_2=5.0 m^2 is the cross-sectional area of the larger cylinder.
Then, from this formula, we can find the force exerted by the piston in the larger cylinder:
F_2=F_1 A_2/A_1 =100 N∙(5.0 m^2)/(0.2 m^2 )=2500 N.